Midland Oil has $1,000 par value bonds outstanding at 8 percent interest. The bonds will mature in 25 years. Use Appendix B and Appendix Dfor an approximate answer but calculate your final answer using the formula and financial calculator methods.

Compute the current price of the bonds if the present yield to maturity is: (Do not round intermediate calculations. Round your final answers to 2 decimal places. Assume interest payments are annual.)
Answer #1

Note: Since no yield to maturity is provided, let me assume it as 10%

Bond Valuation: The value of bond is the present value of the expected cashflows from the bond,discounted at Yield to Maturity(YTM).

Year Cash flow PVAF/PVF@10% Present Value (Cashflow*PVAF/PVF)
1-25 80 9.0770* 726.16
25 1000 0.0923** 92.30

Current Market Price of Bonds =Present Value

= 726.16+92.30

= $818.46

*PVAF = (1-(1+r)^-n)/r

**PVF = 1 / (1+r)^n